OpenStudy (anonymous):

using the addition method,solve each system of equations using the addition method,solve each system of equations x+y=-3 x= 1

OpenStudy (anonymous):

ok for this one all you need to go is plug 1 in for x so you will have 1+y=-3 now subtract the one to the other side and y=-4

OpenStudy (anonymous):

got there now what

OpenStudy (anonymous):

isn't this the substitution method???

OpenStudy (anonymous):

yes now dont i have to find x

OpenStudy (anonymous):

ohh wait we already got x its 1 lol

OpenStudy (anonymous):

no your objective was to find y so y=-4 now you can plug it back into the equation and you get: 1+-4=-3 and that is true

OpenStudy (anonymous):

i got the check done.. can you help me with a few more like these sorta

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

k here it is 3x - 5y=8 5y=10

OpenStudy (anonymous):

3x - 5y=8 5y=10

OpenStudy (anonymous):

Ok so this is how we set the problem up 3x-5y=8 5y=10 now we add downwards, so the y cancels out because -5y+5y= 0 and then we have to add 8+10 which equals 18 so now we have 3x=18 divide 3 by both sides and you are left with x=6 now for the fun part :) since we know x now, we can simply plug it into the top part of the problem 3x-5y=8 plug in 6 for x and we have 3(6)-5y=8 that is really just 18-5y=8 now we subtract 18 from both sides and we have: -5y=-10 if we divide -5 by -10 we get 2 so y=2 plugging all this into the equation 3x-5y=8 3(6)-5(2)= 8 and this is true :)

OpenStudy (anonymous):

thank you so much!.. can you help me with like two more questions like this

OpenStudy (anonymous):

yah do you want me to explain everything again, or just show you the steps quickly? Do you understand how to do it?

OpenStudy (anonymous):

2x+6y=5 2x-3y=-4

OpenStudy (anonymous):

just show me the steps sorta like what u just did but with less words and if i need help ill post heere if i need help

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

2x+6y=5 2x-3y=-4 ok for this problem theres nothing we can make zero right away so we have to multiply by two to get the -3y to -6y so then we can subtract. make sure to multiply by 3 for EVERY variable 2x+6y=5 4x-6y=-8 6x=-3 x=-.5 now we can plug in the x in either equation it doesnt matter 2(-.5)+6y=5 -1+6y=5 6y=6 y=1 check: 4(-.5)-6(1)=8

OpenStudy (anonymous):

sorry thats a -8***

OpenStudy (anonymous):

thanks. i need help with a few more if your okay with helping

OpenStudy (anonymous):

can you try them on your own and show me where you need help?

OpenStudy (anonymous):

wait the equation is not correct cause when u go to check it doesnt fit back in and equal the correct answer

OpenStudy (anonymous):

yes it does 4(-.5)=-2 6(1)=6 -2-6=-8

OpenStudy (anonymous):

oh okay.. can you work through this one with me x+2y=9 3x-2y=7

OpenStudy (anonymous):

can you please try it first on here and show me what you dont understand

OpenStudy (anonymous):

4x=16 x=4 4+2y=9 2y=5 y=5/2

OpenStudy (anonymous):

so u ad 3x +x and got 4x

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

but then you replace 4 from the x to make 4+2y=9

OpenStudy (anonymous):

then what do u do with that 4

OpenStudy (anonymous):

subtract it from 9 because you want the y's on one side

OpenStudy (anonymous):

oohhh

OpenStudy (anonymous):

Does this all make sense, do you have any questions?

OpenStudy (anonymous):

i have one more problem like this but how would u do the check for this one

OpenStudy (anonymous):

x+2y=9 (4)+2(5/2)=9

OpenStudy (anonymous):

wait i got it :) i need help walkin me through one more like u just did

OpenStudy (anonymous):

okay can i see you try it for me please!

OpenStudy (anonymous):

yes and can u help me if i get stuck

OpenStudy (anonymous):

of course, thats our job here at openstudy :)

OpenStudy (anonymous):

x+3y=11 x+3y=7

OpenStudy (anonymous):

u would add 3+3 right to make 6

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

no you cant do that you want to cancel a letter (variable) out so you need to multiply

OpenStudy (anonymous):

grrr

OpenStudy (anonymous):

so what would be the easiest to do

OpenStudy (anonymous):

ugh these are the trickier ones, wait one second well I get my book

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this problem appears like it is a substitution one

OpenStudy (anonymous):

no its an addition method one

OpenStudy (anonymous):

x+3y=11 x+3y=7

OpenStudy (bahrom7893):

do i have to solve that? I can only show the steps

OpenStudy (anonymous):

use addition method

OpenStudy (bahrom7893):

multiply the 2nd equation by -1 and add them

OpenStudy (anonymous):

that would cancel everything out

OpenStudy (bahrom7893):

Actually obviously there are no solutions there.. x+3y cannot equal to 11 and 7 at the same time.

OpenStudy (bahrom7893):

it's inconsistent, or whatever u call it.

OpenStudy (anonymous):

yah that is what i thought

OpenStudy (anonymous):

thanks :)

OpenStudy (shadowfiend):

Kk here we go: x+y=-3 x= 1 You can look at this as: x+y=-3 x+0y= 1 You can then add the negative: x + y = -3 + - (x + 0y = 1) Which is: x + y = -3 + -x = -1 ------------ 0 + y = -4

OpenStudy (shadowfiend):

You already know what x =, because that's part of your two initial equations. Now you know that y = -4, as well, and you've solved the system.

OpenStudy (shadowfiend):

Lol. Never mind, looks like the question evolved from that to something else...

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

no that will not work for x+3y=11

OpenStudy (anonymous):

This problem is inconsistent therefore there is not an anwer that will solve both at the same time

OpenStudy (anonymous):

think about it, its the same variables so they can only equal one number not two at the same time

OpenStudy (anonymous):

okay so it cant be workable..

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

thank you so much for your help..